# CoralVQA: A Large-Scale Visual Question Answering Dataset for Coral Reef Image Understanding

### v1

@@ 1 Introduction
Coral reefs are among the most biodiverse ecosystems, supporting over a quarter of known marine species Knowlton et al. (2010); Fisher et al. (2015). They provide ecological value, support coastal economies through fisheries and tourism, and protect shorelines from storms and erosion Mumby et al. (2008); Bitterwolf et al. (2024). However, human activities and climate change are causing unprecedented global declines Steffen et al. (2007); Hoegh-Guldberg et al. (2019). Effective global coral reef conservation depends on continuous monitoring of benthic communities, with coral reef images serving as a direct and essential source of information. For instance, experts interpret these images to identify species, evaluate their health, and extrapolate findings to downstream systems. However, the capacity to derive insights from coral images is largely restricted to specialists in marine science. Core reef monitoring involves dynamic, multiple tasks–including coral recognition, health diagnostics (e.g., bleaching severity), and habitat quality assessment (e.g., algal symbiosis), all of which demand extensive domain expertise for reliable image interpretation Beijbom et al. (2012a). This barrier restricts the range and variety of problems that can be addressed with coral images (e.g., coral studies in developing countries), and limits the number of potential users (e.g., conservation practitioners, science educators). Therefore, there is a pressing need for new automatic ways that can extract relevant information from coral images without requiring specialized expertise.
@@
The Visual Question Answering (VQA) task–originally developed in computer vision as a user-friendly way for image-based queries–could help bridge this expertise gap Antol et al. (2015); Wu et al. (2017); Kafle, Kanan (2017); Lobry et al. (2020); Zheng et al. (2023). For example, given a coral reef image (the corresponding coral reef image is located in the third row, fourth column in Figure 1) and a question (“Is the coral genus in the upper left corner susceptible to bleaching?”). VQA aims to give the correct answer (“Yes”). This transforms specialized ecological assessment into accessible, high-level semantic information. Recent advances in large vision-and-language models (LVLMs) He et al. (2020); Liu et al. (2023); Bai et al. (2023); Li et al. (2023a); Chen et al. (2024) have demonstrated state-of-the-art performance across various VQA tasks for natural images. However, their extension to the coral reef domain remains limited, primarily due to the absence of comprehensive, high-quality VQA datasets tailored to coral conservation. The construction of coral reef VQA datasets presents substantially greater challenges compared to general-domain VQA, due to two critical factors. (1) Domain-specific annotations. While advances in autonomous underwater vehicles have enabled the large-scale collection of coral reef images, existing annotations are inadequate for supporting VQA dataset development. Key issues include: inconsistent label standards–some corals are labeled by morphology, others by genus, and some at the family level; and inclusion of non-coral categories (e.g., sand and sediments). (2) Multidimensional questions. Current VQA datasets for natural images primarily focus on generic visual concepts (e.g., animals, vehicles) and basic attributes (e.g., quantity, color). This domain-general characteristic makes question-answer pair generation relatively straightforward for non-experts. However, creating coral VQA datasets requires interdisciplinary expertise in both visual content and textual questions. Additionally, to support coral reef conservation, such a dataset must include open-ended question-answer pairs that span multiple ecological and health-related dimensions, such as coral condition, growth status, and symbiotic relationships.
@@
Figure 1: Example of coral reef image and corresponding question-answer pairs across 16 dimensions in CoralVQA dataset. The basic visual attributes of coral reef images are denoted by light blue boxes, while ecological and health-related attributes of coral reef images are indicated by pink boxes.
@@
To address the above challenges, our interdisciplinary team introduces CoralVQA, a large-scale dataset for vision-language understanding of coral reef images, developed through close collaboration with marine biologists possessing extensive expertise in coral conservation. Through data collection, label cleaning and re-annotating, CoralVQA consists of 12,805 high-quality real coral images from multiple marine regions across 3 oceans, including the Atlantic, Indian Ocean, and Pacific Ocean. All coral instances are re-annotated at the genus level, in accordance with the biological taxonomy hierarchy (Kingdom-Phylum-Class-Order-Family-Genus), covering 67 distinct genera from 20 families. CoralVQA encompasses a comprehensive and diverse collection of questions, systematically organized into two key groups: basic visual interpretation; ecological and health-oriented assessment. Each group comprises eight distinct dimensions. We benchmark several state-of-the-art LVLMs and find that CoralVQA presents novel challenges for coral-specific vision–language reasoning. Notably, we observe a significant drop in model performance when answering questions involving images from previously unseen ocean regions or those requiring complex reasoning, such as assessing bleaching coverage. Some examples from CoralVQA are shown in Figure 1. The key contributions of our work are summarized as follows.
@@
  • We design a semi-automatic vision-language coral data collection pipeline, which includes six steps: dataset collection, label cleaning and re-annotating, attribute extraction, question generation, and human verification. Our pipeline can be widely applied to other marine domains, such as visual reasoning tasks in mangrove ecosystem imagery.
@@ 2 Related Works
Coral datasets are fundamental for studying coral reef conservation and assessing coral reef health. Existing works on coral datasets mainly focus on classification or segmentation. For example, Tasmania Coral Point Count (TasCPC) Meyer et al. (2011) contains 1,258 AUV-captured benthic images across 13 underwater object categories (e.g., corals, sand, rock), though coral taxonomic details are unavailable. RSMAS Shihavuddin et al. (2013) is composed of 766 image patches ($\times×256$ pixels) from 8 coral genera, collected by divers from the University of Miami. Benthoz15 Bewley et al. (2015), collected by AUVs from Australian, includes 9,874 labeled images. Despite having 148 categories, only two coral genera follow the Linnaean system. EILAT Shihavuddin (2017) provides 1,123 Red Sea image patches ($\times×64$ pixels),which represent four morphological coral classes, along with categories for favid coral, dead coral, sand, and urchin.
@@
[…]
@@
ATCRC Rashid, Chennu (2020), collected from Curaçao, contains 147 hyperspectral images representing 6 coral genera. HSCR16K Han et al. (2025) contains 16,659 image patches ($\times×224$ pixels) of 10 genera with rich text knowledge. MLC Beijbom et al. (2012b), collected from the island of Moorea, consists of 2,055 coral reef images, including 5 coral genera and 4 non-coral classes. CoralSCOP Zheng et al. (2024), an important step in coral segmentation, contains 41,297 images from sources like YouTube but focuses on segmentation and lacks coral taxonomic details. In this work, we present CoralVQA, a multi-regions, multi-genera, and multi-dimensions VQA dataset, accompanied by comprehensive benchmarks under diverse VQA scenarios.
@@ 3 Pipeline
Figure 2: Pipeline for dataset creation.
@@ 3.1 Dataset Collection
Our CoralVQA collects the real-world coral reef images from 3 sources: MLC dataset Beijbom et al. (2012c), XL Catlin Seaview Survey Project González-Rivero et al. (2019), and the CoralsofttheWorld database (available at coralsoftheworld.org). The XL Catlin Seaview Survey Project (recording the health of coral reefs worldwide), launched in 2012, includes 11,387 images from the Atlantic, Indian, and Pacific Oceans, providing extensive geographic coverage. To enhance both image quantity and taxonomic coverage, we crawled additional 3,420 images from CoralsofttheWorld, a globally recognized source for scientifically validated coral taxonomy and distribution. This resulted in an initial dataset of 16,862 images. After the quality filter using the Underwater Color Image Quality Evaluation metric Yang, Sowmya (2015), the final dataset contains 12,805 high-quality images.
@@ 3.2 Label Cleaning and Re-annotating
Despite the large volume of collected coral reef images, inconsistent annotation standards and non-coral categories pose challenges for direct downstream use. Therefore, we apply a systematic cleaning and re-annotating process guided by the following criteria: 1) removing the coral labels without following the standard coral biological taxonomic system (e.g., Branching Acroporidae and hispidose Acroporidae); 2) eliminating non-coral labels such as sand, fishing gear, and sediments; 3) re-annotating coral instances based on a genus-level coral biological taxonomy hierarchy (from Kingdom and Phylum to Family and Genus and Species). Marine biologists, led by the fourth author, conducted this step. During the process, 50,200 non-coral annotations and location entries were removed, while 97,395 coral-related annotations and their corresponding locations were re-labeled. The final curated dataset contains 12,805 images spanning 20 families and 67 genera.
@@ 3.5 Question-Answer Generation
We utilize the GPT-4o API to automatically generate question-answer pairs. Existing VQA data generation approaches typically rely on feeding textual inputs into large language models. In contrast, we leverage OpenAI’s multimodal capabilities by uploading both textual attributes and coral images through the GPT-4o image API, enabling more authentic and context-aware question–answer generation grounded in visual content. Specifically, based on the image coordinate system, we first divide each image into a 3 $\times$× 3 grid. Using the coral coordinate annotations, we determine the corresponding grid region for each coral instance. The GPT-4o API then automatically assigns directional indicators (e.g., upper left, center, lower right) to support the generation of location-related questions in downstream tasks. To ensure diversity in question generation, we constrain each question type to appear no more than three times per image. To avoid uncertain or vague answers (e.g., uncertain, unknown, or cannot determine), we fix the temperature parameter at 0.3 and incorporate targeted prompt phrases that explicitly require definitive answers. Additionally, we employ the GPT-4o API to perform diversity-oriented paraphrasing, generating questions that differ in linguistic expression while retaining similar semantic intent. We further iteratively refine the prompt templates to better capture domain-specific terminology and improve the overall quality of generated question–answer pairs in the context of coral reef understanding.
@@ 3.6 Human Verification
Although most of the question-answer pairs generated by the GPT-4o API are accurate–due to careful designed prompts–issues still persist. To address these concerns, we perform a three-stage process: 1) manual verification, 2) cross-checking, and 3) expert sampling inspection. In the first stage, twelve students with backgrounds in marine science are recruited to review and correct the question-answer pairs. Corrections are made based on the following criteria: removing hallucinated questions unrelated to the image; fixing incorrect spatial references within questions; revising uncertain and inaccurate answers. After the first stage, most issues are effectively resolved. To further enhance dataset quality, we implement a cross-checking procedure wherein each verified question-answer pair is reviewed by a second annotator. If the accuracy of any subset falls below 95%, it is flagged for re-verification. In the third phase, domain experts perform a quality audit by randomly sampling 10% of the question–answer pairs from each annotator’s contributions. If the sampled accuracy again falls below 95%, the corresponding dataset portion undergoes comprehensive revalidation.
@@ 4.2 Questions-Answers Statistics
CoralVQA contains 277,653 question-answer pairs with 16 question types. The question types in the CoralVQA dataset are categorized into two main groups: (1) basic visual attributes and (2) ecological and health-related attributes. On average, each coral reef image has 21.6 questions. Figure 3(b) illustrates the distribution of question lengths in terms of words, with each question containing an average of 14.4 words. Out of the 16 question types, eight are open-ended, including coral category, color, position, quantity, size, shape, scene type, and health status quantity. These open-ended questions make up 42.1% of all question-answer pairs, totaling 116,960 questions. The diversity of answers to open-ended questions and the variety of visual features in coral reefs present significant challenges for visual language models in coral reef VQA tasks. The remaining question types consist of closed-ended “yes/no” questions, with 73,095 questions having “yes” responses and 87,598 with “no” responses. The distribution of different question types is shown in Figure 4.
Figure 4: Distribution of two question groups.
5 Benchmark VQA Performance
@@ 5.1.1 Models
We evaluate four well-known large vision-language models (LVLMs) methods containing Mini-Gemini(7B) Li et al. (2023b), Qwen2.5VL(7B) Bai et al. (2023), BLIP3 Le Xue (2024), InternVL2.5(8B) Chen et al. (2024), to generate the benchmark results. For the Mini-Gemini, we utilize the OpenCLIP ConvNext Ilharco et al. (2021) as the auxiliary vision encoder and use a 2-layer MLP with GeLU activation for projecting visual features. Following the BLIP3, we utilize Microsoft’s Phi3 Abdin et al. (2024) as the base language model and use Google’s SigLIP-SO400M Zhai et al. (2023) as the vision encoder. Other architectures adhere to the original models. For a fair comparison, we initialize all the models with their own pre-trained weights.

### v2

@@ 1 Introduction
Coral reefs are among the most biodiverse ecosystems, supporting over a quarter of known marine species (Knowlton et al., 2010; Fisher et al., 2015). They provide ecological value, support coastal economies through fisheries and tourism, and protect shorelines from storms and erosion (Mumby et al., 2008; Bitterwolf et al., 2024). However, human activities and climate change are causing unprecedented global declines (Steffen et al., 2007; Hoegh-Guldberg et al., 2019). Effective global coral reef conservation depends on continuous monitoring of benthic communities, with coral reef images serving as a direct and essential source of information. For instance, experts interpret these images to identify species, evaluate their health, and extrapolate findings to downstream systems. However, the capacity to derive insights from coral images is largely restricted to specialists in marine science. Coral reef monitoring involves dynamic, multiple tasks–including coral recognition, health diagnostics (e.g., bleaching severity), and habitat quality assessment (e.g., algal symbiosis), all of which demand extensive domain expertise for reliable image interpretation (Beijbom et al., 2012). This barrier restricts the range and variety of problems that can be addressed with coral images (e.g., coral studies in developing countries), and limits the number of potential users (e.g., conservation practitioners, science educators). Therefore, there is a pressing need for new automatic ways that can extract relevant information from coral images without requiring specialized expertise.
@@
The Visual Question Answering (VQA) task–originally developed in computer vision as a user-friendly way for image-based queries–could help bridge this expertise gap (Antol et al., 2015; Wu et al., 2017; Kafle, Kanan, 2017; Lobry et al., 2020; Zheng et al., 2023; Li et al., 2024). For example, given a coral reef image (the corresponding coral reef image is located in the third row, fourth column in Figure 1) and a question (“Is the coral genus in the upper left corner susceptible to bleaching?”). VQA aims to give the correct answer (“Yes”). This transforms specialized ecological assessment into accessible, high-level semantic information. Recent advances in large vision-and-language models (LVLMs) (He et al., 2020; Liu et al., 2023; Bai et al., 2023; Li et al., 2023a; Chen et al., 2024) have demonstrated state-of-the-art performance across various VQA tasks for natural images. However, their extension to the coral reef domain remains limited, primarily due to the absence of comprehensive, high-quality VQA datasets tailored to coral conservation. The construction of coral reef VQA datasets presents substantially greater challenges compared to general-domain VQA, due to two critical factors. (1) Domain-specific annotations. While advances in autonomous underwater vehicles have enabled the large-scale collection of coral reef images, existing annotations are inadequate for supporting VQA dataset development. Key issues include: inconsistent label standards–some corals are labeled by morphology, others by genus, and some at the family level; and inclusion of non-coral categories (e.g., sand and sediments). (2) Multidimensional questions. Current VQA datasets for natural images primarily focus on generic visual concepts (e.g., animals, vehicles) and basic attributes (e.g., quantity, color). This domain-general characteristic makes question-answer pair generation relatively straightforward for non-experts. However, creating coral VQA datasets requires interdisciplinary expertise in both visual content and textual questions. Additionally, to support coral reef conservation, such a dataset must include open-ended question-answer pairs that span multiple ecological and health-related dimensions, such as coral condition, growth status, and symbiotic relationships.
@@
Figure 1: Example of coral reef image and corresponding question-answer pairs across 16 dimensions in CoralVQA dataset. The basic visual attributes of coral reef images are denoted by light orange boxes, while ecological and health-related attributes of coral reef images are indicated by light green boxes.
@@
To address the above challenges, our interdisciplinary team introduces CoralVQA, a large-scale dataset for vision-language understanding of coral reef images, developed through close collaboration with marine biologists possessing extensive expertise in coral conservation. Through data collection, label cleaning and re-annotating, CoralVQA consists of 12,805 high-quality real coral images from multiple marine regions across 3 oceans, including the Atlantic, Indian Ocean, and Pacific Ocean. All coral instances are re-annotated at the genus level, in accordance with the biological taxonomy hierarchy (Kingdom-Phylum-Class-Order-Family-Genus), covering 67 distinct genera from 20 families. CoralVQA encompasses a comprehensive and diverse collection of questions, systematically organized into two key groups: basic visual interpretation; ecological and health-oriented assessment. Each group comprises eight distinct dimensions. We benchmark several state-of-the-art LVLMs and find that CoralVQA presents novel challenges for coral-specific vision-language reasoning. Notably, we observe a significant drop in model performance when answering questions involving images from previously unseen ocean regions or those requiring complex reasoning, such as assessing bleaching coverage. Some examples from CoralVQA are shown in Figure 1. The key contributions of our work are summarized as follows.
@@
  • We design a semi-automatic vision-language coral data collection pipeline, which includes six steps: dataset collection, label cleaning and re-annotating, attribute extraction, question generation, and human verification. Our pipeline can be widely applied to other marine domains.
@@ 2 Related Works
Coral datasets are fundamental for studying coral reef conservation and assessing coral reef health. Existing works on coral datasets mainly focus on classification or segmentation. For example, Tasmania Coral Point Count (TasCPC) (Meyer et al., 2011) contains 1,258 AUV-captured benthic images across 13 underwater object categories (e.g., corals, sand, rock), though coral taxonomic details are unavailable. RSMAS (Shihavuddin et al., 2013) is composed of 766 image patches ($\times256$ pixels) from 8 coral genera, collected by divers from the University of Miami. Benthoz15 (Bewley et al., 2015), collected by AUVs from Australian, includes 9,874 labeled images. Despite having 148 categories, only two coral genera follow the Linnaean system. EILAT (Shihavuddin, 2017) provides 1,123 Red Sea image patches ($\times64$ pixels), which represent four morphological coral classes, along with categories for favid coral, dead coral, sand, and urchin. ATCRC (Rashid, Chennu, 2020), collected from Curaçao, contains 147 hyperspectral images representing 6 coral genera.
@@
Cls. TasCPC 1,258 * ×
Cls. RSMAS 766 8 ×
Cls. Benthoz15 9,874 2 ×
Cls. EILAT 1,123 * ×
Cls. ATCRC 147 6 ×
Cls. HSCR16K 16,659 10 ×
Seg./Cls. MLC 2,055 5 ×
Seg. CoralSCOP 41,297 * ×
@@
HSCR16K (Han et al., 2025) contains 16,659 image patches ($\times224$ pixels) of 10 genera with rich text knowledge. MLC (Beijbom et al., 2012), collected from the island of Moorea, consists of 2,055 coral reef images, including 5 coral genera and 4 non-coral classes. CoralSCOP (Zheng et al., 2024), an important step in coral segmentation, contains 41,297 images from sources like YouTube but focuses on segmentation and lacks coral taxonomic details. In this work, we present CoralVQA, a multi-regions, multi-genera, and multi-dimensions VQA dataset, accompanied by comprehensive benchmarks under diverse VQA scenarios.
@@ 3 Pipeline
Figure 2: Pipeline for dataset creation. CoralVQA is developed through a six-stage semi-automated vision-language workflow: dataset collection, label cleaning and re-annotating, attribute extraction, prompt engineering, question generation, and human verification.
@@ 3.1 Dataset Collection
Our CoralVQA collects the real-world coral reef images from 3 sources: MLC dataset (Beijbom et al., 2012), XL Catlin Seaview Survey Project (González-Rivero et al., 2019), and the CoralsofttheWorld database (available at coralsoftheworld.org). The XL Catlin Seaview Survey Project (recording the health of coral reefs worldwide), launched in 2012, includes 11,387 images from the Atlantic, Indian, and Pacific Oceans, providing extensive geographic coverage. To enhance both image quantity and taxonomic coverage, we crawled additional 3,420 images from CoralsofttheWorld, a globally recognized source for scientifically validated coral taxonomy and distribution. This resulted in an initial dataset of 16,862 images. After the quality filter using the Underwater Color Image Quality Evaluation metric (Yang, Sowmya, 2015), the final dataset contains 12,805 high-quality images.
@@ 3.2 Label Cleaning and Re-annotating
Despite the large volume of collected coral reef images, inconsistent annotation standards and non-coral categories pose challenges for direct downstream use. Therefore, we apply a systematic cleaning and re-annotating process guided by the following criteria: 1) removing the coral labels without following the standard coral biological taxonomic system (e.g., branching Acroporidae and hispidose Acroporidae); 2) eliminating non-coral labels such as sand, fishing gear, and sediments; 3) re-annotating coral instances based on a genus-level coral biological taxonomy hierarchy (from Kingdom and Phylum to Family and Genus and Species). Marine biologists, led by the fourth author, conducted this step. During the process, 50,200 non-coral annotations and location entries were removed, while 97,395 coral-related annotations and their corresponding locations were re-labeled. In this step, the major issue is inconsistent annotation standards and incorrect coral annotations. To address the problem, we have every image re-examined by three marine scientists after the label cleaning and re-annotating step, and adopt a majority voting approach to determine the final annotations. In fact, only a very small set of images (203 out of 12,805, <2 %) contains errors. The final curated dataset contains 12,805 images spanning 20 families and 67 genera.
@@ 3.5 Question-Answer Generation
We utilize the GPT-4o API to automatically generate question-answer pairs. Existing VQA data generation approaches typically rely on feeding textual inputs into large language models. In contrast, we leverage OpenAI’s multimodal capabilities by uploading both textual attributes and coral images through the GPT-4o image API, enabling more authentic and context-aware question–answer generation grounded in visual content. Specifically, based on the image coordinate system, we first divide each image into a 3 $\times$ 3 grid. Using the coral coordinate annotations, we determine the corresponding grid region for each coral instance. The GPT-4o API then automatically assigns directional indicators (e.g., upper left, center, lower right) to support the generation of location-related questions in downstream tasks. To ensure diversity in question generation, we constrain each question type to appear no more than three times per image. Additionally, we employ the GPT-4o API to perform diversity-oriented paraphrasing, generating questions that differ in linguistic expression while retaining similar semantic intent. We further iteratively refine the prompt templates to better capture domain-specific terminology and improve the overall quality of generated question–answer pairs in the context of coral reef understanding. In this step, the major issue is about the reproducibility issues of GPT-4o generated question-answer pairs. To address the issue, we set the lower temperature parameter of 0.3 and incorporate targeted prompt phrases that explicitly require definitive answers. We randomly sampled 1,000 questions, with each question being answered 10 times by GPT-4o. The average number of consistent answers was 8.2 times. Although GPT-4o generated answers exhibit minor inconsistencies, since all question answers undergo multiple rounds of manual verification, the impact of answer inconsistencies is further mitigated.
@@ 3.6 Human Verification
Even with carefully designed prompts, errors persist in the GPT-4o API’s question–answer pairs. To address these concerns, we perform a three-stage process: 1) manual verification, 2) cross-checking, and 3) expert sampling inspection. In the first stage, twelve students with backgrounds in marine science are recruited to verify the coral image and question-answer pairs manually and fix the errors resulted from using GPT-4o. Each of them needed to deal with roughly 1,000 images. In total, there were 12 subsets of images and corresponding question-answer pairs. Corrections are made based on the following criteria: removing hallucinated questions unrelated to the image; fixing incorrect spatial references within questions; revising uncertain and inaccurate answers. After the first stage, most issues are effectively resolved. To further enhance dataset quality, we implement a cross-checking procedure wherein each verified question-answer pair is reviewed by a second annotator. Each student cross-checked the subset of images verfied by another student. If the consistency of cross-checking is lower than 95% in any specific subset, a third inspector was assigned to check these inconsistencies again and perform necessary fixes. In the third phase, domain experts inspected 10% images and corresponding question-answer pairs randomly choosen from each subset. If the accuracy of a subset’s sample was lower than 95% (judged by the assigned expert), that subset had to be re-verified again. We compiled statistics on the manual pruning of GPT-4o responses. During the human verification stage, we modified 13.4% of the questions generated by GPT-4o and 56.8% of the answers generated by GPT-4o required manual pruning.
@@ 4.2 Questions-Answers Statistics
CoralVQA contains 277,653 question-answer pairs with 16 question types. The question types in the dataset are categorized into two main groups: (1) basic visual attributes and (2) ecological and health-related attributes. On average, each coral reef image has 21.6 questions. Figure 3(b) illustrates the distribution of question lengths in terms of words, with each question containing an average of 14.4 words. Out of the 16 question types, eight are open-ended, including coral category, color, position, quantity, size, shape, scene type, and health status quantity. These open-ended questions make up 42.1% of all question-answer pairs, totaling 116,960 questions. The diversity of answers to open-ended questions and the variety of visual features in coral reefs present significant challenges for visual language models in coral reef VQA tasks. The remaining question types consist of closed-ended “yes/no” questions, with 73,095 questions having “yes” responses and 87,598 with “no” responses. The distribution of different question types is shown in Figure 4.
5 VQA Benchmark Performances
@@ 5.1.1 Models
To establish benchmark results, we conduct evaluations on six representative large vision-language models (LVLMs). Among them, the open-source models are Mini-Gemini(7B) (Li et al., 2023b), Qwen2.5VL(7B) (Bai et al., 2023), BLIP3 (Le Xue, 2024), and InternVL2.5(8B) (Chen et al., 2024). In addition, we also include two leading closed-source models, GPT-4o (Hurst et al., 2024) and Claude3.5 Haiku (Anthropic, 2024). For the Mini-Gemini, we utilize the OpenCLIP ConvNext (Ilharco et al., 2021) as the auxiliary vision encoder and use a 2-layer MLP with GeLU activation for projecting visual features. Following the BLIP3, we utilize Microsoft’s Phi3 (Abdin et al., 2024) as the base language model and use Google’s SigLIP-SO400M (Zhai et al., 2023) as the vision encoder. Other architectures adhere to the original models. For a fair comparison, we initialize all open-source models with their own pre-trained weights. Additional training details are provided in the appendix.